OpenStudy (anonymous):

Evaluate the function rule to find the range of each function for the domain {-3, 0, 5}. f(x) = x² − 5 A. {-14, .5, 10} B. {-14, -5, 20} C. {-5, 4, 20} D. {4, 5, 10}

OpenStudy (texaschic101):

domain is all the x values. so lets start subbing you numbers in for x.. f(x) = x^2 - 5 subbing in -3 for x f(-3) = -3^2 - 5 f(-3) = 9 - 5 f(-3) = 4 subbing in 0 for x f(0) = 0^2 - 5 f(0) = -5 subbing in 5 for x f(5) = 5^2 - 5 f(5) = 25 - 5 f(5) = 20 so your range, the y values, are : -5,4,20

OpenStudy (anonymous):

tysm! c: